OpenStudy (anonymous):

Rewrite the subtraction problem as an addition problem. 172 – x = 292

OpenStudy (anonymous):

Which of the following represents the largest number? A) 21 base eight B) 18 base nine C) 30 base four D) 18 base eleven

OpenStudy (pottersheep):

Sorry slight mistake 172 – x = 292 -x = 292 - 172 -x = -120 x = 120

OpenStudy (jhannybean):

172 - x = 292 subtraction problem because of the " - x" 172 - (-x) = 292 We can change this into an addition problem and rewrite it as 172 + x = 292 Now we solve for x by subtracting 172 from both sides 172 -172 + x = 292 -172 Therefore x = 292 - 172 = 120

OpenStudy (anonymous):

to rewrite this as a subtraction problem you must add an x to both sides 172-x=292 +x +x the x on the left side would cancel out and the problem would be rewritten like this 172=x+292 because you dont have to solve for x you only need to rewrite it as an addition problem. a problem isnt solve yet :D hope you understand

OpenStudy (anonymous):

but f you do have to solve you isolate the x